Dragons18:

A geologist and South America discovers a bird it’s cold in the content 77% of its original amount of C – 14. On the age of the bird skeletons the nearest year. A geologist in South America discovers a bird it’s cold in the content 77% of its original amount of C – 14. Find the age of the bird skeleton of the nearest year.

sillybilly123:

The half-life of C-14 is roughly 5,730 years So: \(A(t) = A_o \left( \dfrac{1}{2}\right)^{\frac{t}{5,730}}\) Where ... \(A\) is a proxy for any metric: mass, weight, #, etc Your question is garbled. But birds are really just the remnants of the flying dinosaurs so I am guessing it is a real question. There is absolutely no evidence for dragons, though :(